Find the length of a rectangular piece of fiberglass if its length is 3 times the width and the area is 591.Round to the nearest inch.
Answer by checkley77(12844)

L=3W
LW=591
3W*W=591
3W^2=591
W^2=591/3
W^2=197
W=SQRT197
W=14 INCHES.
L=3*14=42 INCHES.
PROOF:
14*42=591
588~591